OpenStudy (anonymous):

Please help me understand!! Solve 4w − 7k = 28 for k. Show your work

OpenStudy (anonymous):

@peachpi last one, I promise.. sorry :)

OpenStudy (anonymous):

what do you have so far?

OpenStudy (anonymous):

+4w to both sides, to get 32w on the right side? so 7k=32w?

OpenStudy (anonymous):

4w is positive on the left side, so you need to subtract it. When you subtract, 32 and 4w are not like terms, so they DO NOT COMBINE. -7k = 32 - 4w What comes next?

OpenStudy (anonymous):

divide by -7? i mean 32 divided by -7

OpenStudy (anonymous):

wait hold up. it was 28 on the right side, so -7k = 28 - 4w then yes, divide by -7

OpenStudy (anonymous):

so k=-4-4w?

OpenStudy (anonymous):

You have to divide the whole side by -7. Both of these are correct